### v1

# Existence results for a biofilm free-boundary problem with dominant detachment

Abstract
This work addresses the existene and uniqueness of a Wanner-Gujer free boundary problem modeling biofilms with a prevailing detachment. This result extends the some works from tumor gowth modeling as well as those from the biofilm modeling field. Besides the local existence and uniqueness, the continuous dependence on initial and boundary data and the global existence is also given. The local existence is obtained by using fixed point arguments combined with semigroup theory while the global existence is deduced using invariance regions and energy estimates.
keywords:
Biofilms , Free boundary problem , Semigroup , Anisotropic spaces , Hyperbolic and Parabolic PDE , Energy estimates MSC 35A20 , 35A08 , 35A07 , 35A22 , 35C15 , 35R35
journal: Applied Mathematics letters
[…]
organization=Départment de Mathématique-Physique, Institut Supérieur Pédagogique de Bukavu,addressline=30 Avenue Kibombo, city=Bukavu, postcode=854, country=Democratic Republic of Congo
1 Introduction
We consider the following free bounddary prroblem which is known as Wanner-Gujer model:
$\partial_{t}X_{i}+u\partial_{x}X_{i}=f_{i}(\boldsymbol{X},\boldsymbol{S}),~0\leq x\leq L(t),~t>0,~i=1,\cdots,n;$ (1a)
$X_{i}(x,0)=\varphi_{i}(x),~0\leq x\leq L_{0},~~i=1,\cdots,n~;$ (1b)
$\partial_{x}u=g(\boldsymbol{X},\boldsymbol{S}),~~0<x\leq L(t);$ (1c)
u(0,t)=0,t>0.u(0,t)=0,~~t>0. (1d)
$\partial_{t}S_{j}-D_{j}\partial_{xx}S_{j}=h_{j}(\boldsymbol{X},\boldsymbol{S}),~0<x<L(t),~t>0,~j=1\cdots,m;$ (2a)
$S_{j}(x,0)=\theta_{j}(x),~t>0,~j=1\cdots,m;$ (2b)
$\partial_{x}S_{j}(0,t)=0;~~S_{j}(L(t),t)=\psi_{j}(t),~t>0;~j=1,\cdots,m.$ (2c)
$\dot{L}(t)=u(L(t),t)-\lambda L^{2}(t),~\lambda>0,~t>0;$ (3a)
$L(0)=L_{0}.$ (3b)
This model describes the spatio-temporal evolution of biofilm components. Each biomass species $X_{i}$ grows or decays according to $f_{i}$, and is transported by the velocity field $u$. The substrates $S_{j}$ diffuse and are consumed by the biomass through the terms $h_{j}$. In general, $f_{i}$ and $h_{j}$ follow Monod-type kinetics [1, 2, 3, 4]. The biofilm thickness $L(t)$ evolves over time as the biofilm grows and detaches, with the detachment rate proportional to $L^{2}(t)$. The following notations $\boldsymbol{X}=(X_{1}\cdots,X_{n})$ and $\boldsymbol{S}=(S_{1},\cdots,S_{m})$ will be used from now on. More details about the derivation and other comments can be found in [5, 6, 7, 8, 9]. The mathematical investigation of its qualitative properties is less expanded compared to the amount of the works applying it to modeling different process in biotechnology and wastewater treatment [10, 11, 12].
When $\lambda=0$ the problem can be treated as the Stefan problem or as tumor growth modeling problems [13, 14, 15, 16, 17]. In this paper we consider the case where $\lambda\neq 0$ which is general and frequently used in wastewater treatment and biotechnology engineering. Indeed, while such models are classical in environmental and bioprocess engineering, their analytical well-posedness has received comparatively little attention except in [18, 19] where it was studied by using method of characteristics. But this approach does not allow to address the global existence due to potential blow-up. Hence, this work addresses the question of local and global existence and uniqueness by using a different technique. More precisely, it establishes a complete local well-posedness theory for the coupled hyperbolic–parabolic free-boundary system, formulated after a nonlinear rescaling to a fixed reference domain. The proof combines maximal $L^{p}$-regularity for parabolic systems with a fixed-point argument coupling the transport and diffusion equations through the moving boundary ODE.
We unify the analysis under minimal regularity assumptions: the results hold for all $p>\tfrac{3}{2}$ in one spatial dimension, which is sharp for the embeddings needed to control the nonlinear couplings. This extends some works from tumor growth modeling that required stronger assumptions for $\lambda=0.$ For example, it was required that $p>5/2$ in [15] and $p>5$ in [20] while a classical $C^{2}$ and Hölder $C^{\alpha+2,\alpha+\frac{1}{2}}$ regularity was respectively obtained in [21] and [22].
The approach used yields not only local existence and uniqueness but also continuous dependence on data, positivity preservation of biomass and substrates, and a continuation criterion that allows extension to global-in-time solutions under dissipative (detachment-dominated) conditions.
Finally, the analysis proposed covers both the parabolic and hyperbolic subproblems as separate lemmas (Lemmas 3.13.2), providing reusable well-posedness results for related coupled reaction–transport systems beyond biofilm modeling.
These results therefore bridge a gap between applied biofilm models widely used in wastewater and bioengineering applications and their mathematical analysis, offering an adapted existence and uniqueness theory consistent with the nonlinear coupling and free-boundary dynamics characteristic of the Wanner–Gujer modeling.
We introduce the notations which will be used in this work: Throughout the paper let $p>\tfrac{3}{2}$ be fixed. For any open interval $\Omega\subset\mathbb{R}$ and integer $k\geq 1$, we denote by $W^{k,p}(\Omega)$ the usual Sobolev space, with norm
$\|u\|_{W^{k,p}(\Omega)}=\sum_{|\alpha|\leq k}\|\partial_{x}^{\alpha}u\|_{L^{p}(\Omega)}$.
For a positive continuous function $L=L(t)$ defined on $0,T_{1}]$ and for some $T_{1}>0$, we set
$\Omega_{L}:=\{(x,t)\in\mathbb{R}\times(0,T_{1}):0<x<L(t)\}\quad\text{and }\quad\overline{\Omega}_{L}\text{ its closure.}$
For functions defined on $\Omega_{L}$, we use the standard parabolic Sobolev space
$W^{2,1}_{p}(\Omega_{L})=\bigl\{u\in L^{p}(\Omega_{L}):\partial_{t}u,\;\partial_{x}u,\;\partial_{xx}u\in L^{p}(\Omega_{L})\bigr\},$
endowed with the norm
$\|u\|_{W^{2,1}_{p}(\Omega_{L})}:=\|u\|_{L^{p}(\Omega_{L})}+\|\partial_{t}u\|_{L^{p}(\Omega_{L})}+\|\partial_{x}u\|_{L^{p}(\Omega_{L})}+\|\partial_{xx}u\|_{L^{p}(\Omega_{L})}.$
Let $D_{p}(\Omega)$ denote the trace of $W^{2,1}_{p}(\Omega\times(0,T_{1}))$ at time $t=0$:
$D_{p}(\Omega)=\{\,\theta\in L^{p}(\Omega):\exists\,u\in W^{2,1}_{p}(\Omega\times(0,T_{1})),\;u(\cdot,0)=\theta\,\}.$
The trace theorem (see for example [23, 24, 25]) implies that for $p>\tfrac{3}{2}$,
$W^{2,1}_{p}(\Omega\times(0,T_{1}))\hookrightarrow C([0,T_{1}];W^{2-2/p,p}(\Omega))\hookrightarrow C(\overline{\Omega}\times[0,T_{1}]),$
so that time traces and boundary values are well defined and continuous. The norm in $D_{p}(\Omega)$ is defined by
$\|\theta\|_{D_{p}(\Omega)}=\inf\bigl\{T_{1}^{-1/p}\|u\|_{W^{2,1}_{p}(\Omega\times(0,T_{1}))}:u\in W^{2,1}_{p}(\Omega\times(0,T_{1})),\ u(\cdot,0)=\theta\bigr\}.$
In particular, if $\theta\in W^{2-2/p,p}(\Omega)$, then $\theta\in D_{p}(\Omega)$ and $\|\theta\|_{D_{p}(\Omega)}\leq\|\theta\|_{W^{2-2/p,p}(\Omega)}$. We will consider $\Omega=(0,1)$ and $\overline{\Omega}=[0,1]$ (see section 2). Since $p>\tfrac{3}{2}$, we have the continuous embedding $W^{2,1}_{p}((0,1)\times(0,T_{1}))\hookrightarrow C([0,1]\times[0,T_{1}])$, which is the regularity level used in Lemmas 3.13.2, and Theorems 3.34.1.
Hence all subsequent results are formulated for $p>\tfrac{3}{2}$, which is the minimal condition ensuring the required embeddings and traces in one spatial dimension.
2 Reformualtion of the problem
In order to investigate the existence and uniqueness of the problem (1)–(3) we make the following change of variables :
$z=\frac{x}{L(t)},~\tilde{t}=\int_{0}^{t}\frac{d\tau}{L(\tau)},~R(\tilde{t})=L(t),~C_{j}(z,\tilde{t})=S_{j}(x,t),j=1,\cdots,m;$
$Y_{i}(z,\tilde{t})=X_{i}(x,t),i=1,\cdots,n;~~v(x,\tilde{t})=L(t)u(x,t).$
To simplify the notations we omit the tilde on $t,$ hence we get the following intial boundary value problem
$\partial_{t}Y_{i}-zv(1,t)\partial_{z}Y_{i}=R^{2}(t)f_{i}(\boldsymbol{Y},\boldsymbol{C})=:F_{i}(\boldsymbol{Y},\boldsymbol{C}),0\leq z\leq 1,~t>0,~i=1,\cdots,n;$ (4a)
$Y_{i}(z,0)=\varphi_{i}(z),~0\leq z\leq 1,~~i=1,\cdots,n~;$ (4b)
$v(z,t)=u(z,t)-zu(1,t)\qquad\text{for}\quad 0\leq z\leq 1,\quad t>0;$ (4c)
$\begin{array}[]{r}\partial_{t}C_{j}-zv(1,t)\partial_{z}C_{j}(z,t)-D_{j}\partial_{zz}C_{j}=R^{2}(t)h_{j}(\boldsymbol{Y},\boldsymbol{C})=:H_{j}(\boldsymbol{Y},\boldsymbol{C}),\\ 0<z<1,~t>0,~j=1,\cdots,m;\end{array}$ (5a)
$C_{j}(z,0)=\theta_{j}(z),~t>0,~j=1,\cdots,m;$ (5b)
$\partial_{z}C_{j}(0,t)=0;~~C_{j}(1,t)=\psi_{j}(t),~t>0;~j=1,\cdots,m.$ (5c)
$v(z,t)=R^{2}(t)\int_{0}^{z}g(\boldsymbol{Y}(\xi,t),\boldsymbol{C}(\xi,t))d\xi$ (6a)
$v(1,t)=R^{2}(t)\int_{0}^{1}g(\boldsymbol{Y}(\xi,t),\boldsymbol{C}(\xi,t))d\xi$ (6b)
$\dot{R}(t)=R^{2}(t)v(1,t)-\lambda R^{4}(t),~t>0;$ (7a)
$R(0)=R_{0}.$ (7b)
Define the following vector-valued functions : $\boldsymbol{C}=(C_{1},\cdots,C_{m})^{T},\boldsymbol{Y}=(Y_{1},\cdots,Y_{n})^{T},$ $\boldsymbol{F}=(F_{1},\cdots,F_{n})^{T},\boldsymbol{H}=(H_{1},\cdots,H_{m})^{T},\boldsymbol{B}(t)=B(t)\boldsymbol{I}_{m}$ where $B(t)=zv(1,t)$ and $\boldsymbol{I}_{m}$ the identity matrix of order $m.$ This system is equivalent to equations (1)–(3) and will be the focus of the rest of the work. We will first give primary lemmas which will be used in the proof of the main results.
3 Primary lemmas and local existence
With the notations introduced in the previous section we recast the parabolic system (5a) into a vector form as follows
$\left\{\begin{array}[]{r}\partial_{t}\boldsymbol{C}(z,t)=\partial_{zz}\boldsymbol{C}(z,t)+\boldsymbol{B}(t)\partial_{z}\boldsymbol{C}(z,t)+\boldsymbol{H}(\boldsymbol{Y}(z,t),\boldsymbol{C}(z,t)),\\ ~(x,t)\in\Omega\times[0,T_{1}];\\ \\ \boldsymbol{C}(z,0)=\boldsymbol{\theta}(z),~z\in\overline{\Omega};\\ \\ \partial_{z}\boldsymbol{C}(0,t)=0;~~\boldsymbol{C}(1,t)~=\boldsymbol{\Psi}(t),~t>0.\end{array}\right$. (8)
where $\Omega=(0,1)$ and $\overline{\Omega}=[0,1].$
We make the following assumptions
  • (i)
    $\boldsymbol{B}\in C^{\alpha}([0,T_{1}];\mathbb{R}^{m\times m})$ for some $\alpha\in(0,1];$
  • (ii)
    $\boldsymbol{H}:\mathbb{R}^{m}\times\mathbb{R}^{m}\to\mathbb{R}^{m}$ is globally Lipschitz with constant $L_{H};$
  • (iii)
    $\boldsymbol{Y}\in L^{\infty}(\Omega\times(0,T_{1}))^{m};$
  • (iv)
    $\boldsymbol{\theta}\in W^{2-2/p,p}(\Omega)^{m}$ for some $p\in(\frac{3}{2},\infty);$
  • (v)
    $\boldsymbol{\Psi}\in W^{1-\frac{1}{2p},p}(0,T_{1})^{m}$ with $\boldsymbol{\theta}(1)=\boldsymbol{\Psi}(0);$
  • (vi)
    (Second-order compatibility) $\partial_{zz}\boldsymbol{\theta}(1)+\boldsymbol{B}(0)\partial_{z}\theta(1)+\boldsymbol{H}(\boldsymbol{Y}(1,0),\boldsymbol{\theta}(1))=\frac{d}{dt}\boldsymbol{\Psi}(0).$
Lemma 3.1
Under the assumptions (i)–(vi) there exists $T^{*}\in(0,T_{1}]$ and a unique solution $\boldsymbol{C}\in W^{2,1}_{p}(\Omega\times(0,T^{*}))^{m}$ to (8). Moreover the following estimate holds
$\|\boldsymbol{C}\|_{W^{2,1}_{p}(\Omega\times(0,T^{*}))^{m}}\leq M\left(\|\boldsymbol{\theta}\|_{W^{2-2/p,p}(\Omega)^{m}}+\|\boldsymbol{H}(\boldsymbol{Y},0)\|_{L^{p}(\Omega\times(0,T^{*}))^{m}}\right$.
$\left.+\|\boldsymbol{\Psi}\|_{W^{1-\frac{1}{2p},p}(0,T^{*})^{m}}+1\right)$ (9)
where $M>0$ depends on $p,T^{*},\|\boldsymbol{B}\|_{C^{\alpha}([0,T^{*}])}$, and $L_{H}$.
Proof. The proof is divided into several steps, the first of which is the parabolic lifiting of boundary data. In fact, since $\boldsymbol{\Psi}\in W^{1-\frac{1}{2p},p}(0,T_{1})^{m}$, by the parabolic trace theorem (see for example [26, 27]), there exists $\tilde{\boldsymbol{\Psi}}\in W^{2,1}_{p}(\Omega\times(0,T_{1}))^{m}$ such that:
  • 1.
    $\tilde{\boldsymbol{\Psi}}(1,t)=\boldsymbol{\Psi}(t)$ for $t\in(0,T_{1});$
  • 2.
    $\tilde{\boldsymbol{\Psi}}(z,0)=\boldsymbol{\theta}(z)$ near $z=1$ (can be arranged via localization);
  • 3.
    $\|\tilde{\boldsymbol{\Psi}}\|_{W^{2,1}_{p}(\Omega\times(0,T_{1}))^{m}}\leq c\left(\|\boldsymbol{\Psi}\|_{W^{1-\frac{1}{2p},p}(0,T_{1})^{m}}+\|\boldsymbol{\theta}\|_{W^{2-2/p,p}(\Omega)^{m}}\right).$
These conditions will later allow to apply maximal regularity in this proof.
The second step is the homogenization and abstract formulation by which we will convert the problem into an abstract evolution equation which later will be used in the next steps.
Let us define $\tilde{\boldsymbol{C}}(z,t)=\boldsymbol{C}(z,t)-\tilde{\boldsymbol{\Psi}}(z,t)$. Then $\tilde{\boldsymbol{C}}$ satisfies:
$\begin{cases}\partial_{t}\tilde{\boldsymbol{C}}=\partial_{zz}\tilde{\boldsymbol{C}}+\boldsymbol{B}(t)\partial_{z}\tilde{\boldsymbol{C}}+\tilde{\boldsymbol{H}}(t),&(z,t)\in\Omega\times(0,T_{1}]\\ \partial_{z}\tilde{\boldsymbol{C}}(0,t)=0,\quad\tilde{\boldsymbol{C}}(1,t)=0,&t\in(0,T_{1}]\\ \tilde{\boldsymbol{C}}(z,0)=\tilde{\boldsymbol{C}}_{0}(z):=\boldsymbol{C}_{0}(z)-\tilde{\boldsymbol{\Psi}}(z,0),&z\in\Omega\end{cases}$ (10)
where
$\tilde{\boldsymbol{H}}(z,t)=\boldsymbol{H}(\boldsymbol{Y}(z,t),\tilde{\boldsymbol{C}}(z,t)+\tilde{\boldsymbol{\Psi}}(z,t))+\boldsymbol{B}(t)\partial_{z}\tilde{\boldsymbol{\Psi}}+\partial_{zz}\tilde{\boldsymbol{\Psi}}-\partial_{t}\tilde{\boldsymbol{\Psi}}$.
On $\mathbb{X}=L^{p}(\Omega)^{m}$, we define the following differential operator
$\mathcal{A}(t)\phi=(\partial_{zz}+\boldsymbol{B}(t)\partial_{z})\phi$
on the domain
$D(\mathcal{A}(t))=D:=\{\phi\in W^{2,p}(\Omega)^{m}:\partial_{z}\phi(0)=0,\phi(1)=0\}.$
The problem becomes:
$\begin{cases}\frac{d\tilde{\boldsymbol{C}}}{dt}(t)=\mathcal{A}(t)\tilde{\boldsymbol{C}}(t)+\mathcal{F}(t,\tilde{\boldsymbol{C}}(t)),&t\in(0,T^{*}]\\ \tilde{\boldsymbol{C}}(0)=\boldsymbol{\theta}\end{cases}$ (11)
with $\mathcal{F}(t,\phi)(z)=\boldsymbol{H}(\boldsymbol{Y}(z,t),\phi(z)+\tilde{\boldsymbol{\Psi}}(z,t))+(\boldsymbol{B}(t)\partial_{z}+\partial_{zz}-\partial_{t})\tilde{\boldsymbol{\Psi}}(z,t)$. Let $\mathcal{A}_{0}\phi=\partial_{zz}\phi$ with domain $D$. On one hand, by [28, Theorem 2.7, Chap.7] and [29, Chap. 8-9], $\mathcal{A}_{0}$ is strongly elliptic and generates an analytic semigroup on $\mathbb{X}$.
On the otehr hand, the perturbation $P(t)\phi=\boldsymbol{B}(t)(z\partial_{z}\phi)$ is bounded using the one-dimensional Gagliardo-Nirenberg inequality [30, 31]:
$\|\partial_{z}\phi\|_{L^{p}}\leq C\|\phi\|_{W^{2,p}}^{1/2}\|\phi\|_{L^{p}}^{1/2}\leq\varepsilon\|\phi\|_{W^{2,p}}+C_{\varepsilon}\|\phi\|_{L^{p}}$
we obtain:
$\|P(t)\phi\|_{L^{p}}\leq\|\boldsymbol{B}\|_{L^{\infty}}\left(\varepsilon\|\mathcal{A}_{0}\phi\|_{L^{p}}+C_{\varepsilon}\|\phi\|_{L^{p}}\right)$
Thus $P(t)$ is $\mathcal{A}_{0}$-bounded with relative bound 0 and $\mathcal{A}(t)$ generates an analytic semigroup. Since $t\mapsto\boldsymbol{B}(t)$ is Hölder continuous and $D$ is constant, it follows that the Acquistapace-Terreni conditions [32] are satisfied, yielding an evolution system $\{U(t,s)\}_{0\leq s\leq t\leq T_{1}}$. To prove the existence of a unique fixed-point we investigate the maximal regularity. Let us define the Banach space:
$\mathbb{E}_{T^{*}}=W^{1,p}(0,T^{*};X)\cap L^{p}(0,T^{*};D)$
which is isomorphic to $W^{2,1}_{p}(I\times(0,T^{*}))^{m}$ by standard parabolic theory. We also need to the following closed ball
$\mathcal{B}_{T^{*}}=\left\{\phi\in\mathbb{E}_{T^{*}}:\phi(0)=\tilde{\boldsymbol{C}}_{0},\ \|\phi\|_{\mathbb{E}_{T^{*}}}\leq R^{*}\right\}$
for some $R^{*}>0.$ With these tools at hand we can rewrite the abstract evolution equation (11) into an integal form as follows
$\Phi(\phi)(t)=U(t,0)\boldsymbol{\theta}+\int_{0}^{t}U(t,s)\mathcal{F}(s,\phi(s))\,ds$ (12)
The last part of the proof is dedicated to show that the right hand side of the integral equation (12) has a unique fixed point. The following standard arguments are used: first it is easy to show that [33]:
$\|\Phi(\phi)\|_{\mathbb{E}_{T^{*}}}\leq M_{1}\left(\|\boldsymbol{\theta}\|_{W^{2-2/p,p}}+\|\mathcal{F}(\cdot,\phi)\|_{L^{p}(0,T^{*};\mathbb{X})}\right).$
In particular the the nonlinear term satisify
$\|\mathcal{F}(\cdot,\phi)\|_{L^{p}(0,T^{*};\mathbb{X})}$ $\leq L_{H}\left(\|\phi\|_{L^{p}(0,T^{*};\mathbb{X})}+\|\tilde{\boldsymbol{\Psi}}\|_{L^{p}(0,T^{*};\mathbb{X})}\right)$
$\quad+\|\boldsymbol{H}(\boldsymbol{Y},0)\|_{L^{p}(0,T^{*};\mathbb{X})}+C\|\tilde{\boldsymbol{\Psi}}\|_{W^{2,1}_{p}}$
By the embedding $\mathbb{E}_{T^{*}}\hookrightarrow L^{\infty}(0,T^{*};W^{2-2/p,p}(\Omega))$ [34] we have
$\|\phi\|_{L^{p}(0,T^{*};\mathbb{X})}\leq(T^{*})^{1/p}\|\phi\|_{L^{\infty}(0,T^{*};\mathbb{X})}\leq C(T^{*})^{1/p}\|\phi\|_{\mathbb{E}_{T^{*}}}$
Using the assumption (ii) we get the following Lipschitz estimate
$\|\mathcal{F}(\cdot,\phi_{1})-\mathcal{F}(\cdot,\phi_{2})\|_{L^{p}(0,T^{*};\mathbb{X})}\leq L_{H}C(T^{*})^{1/p}\|\phi_{1}-\phi_{2}\|_{\mathbb{E}_{T^{*}}}.$
We therefore choose $R^{*}$ large enough and $T^{*}$ small enough so that:
$M_{1}\left(\|\tilde{\boldsymbol{C}}_{0}\|_{W^{2-2/p,p}}+L_{H}(R^{*}+\|\tilde{\boldsymbol{\Psi}}\|)+\|\boldsymbol{H}(\boldsymbol{Y},0)\|+C\|\tilde{\boldsymbol{\Psi}}\|\right)\leq R^{*}$
and
$M_{1}L_{H}C(T^{*})^{1/p}<1$
Hence $\Phi$ is a contraction on $\mathcal{B}_{T^{*}}$. By standard bootstrapping and uniqueness of mild/strong solutions, this solution coincides with the unique classical solution in $W_{p}^{2,1}$
The fixed point $\tilde{\boldsymbol{C}}\in\mathbb{E}_{T^{*}}$ gives the solution to the homogeneous problem. Returning to $\boldsymbol{C}=\tilde{\boldsymbol{C}}+\tilde{\boldsymbol{\Psi}}$, we obtain the solution to the original problem.
The estimate follows from the fixed point property and the linear estimates. The second-order compatibility condition ensures that the solution maintains the $W^{2,1}_{p}$ regularity up to $t=0$. $\square$
Lemma 3.2
[Well-posedness and estimates for a transport–reaction system] Let $v\in C^{1}(\overline{\Omega}\times[0,T_{1}])$ satisfy $\|v\|_{L^{\infty}}+\|\partial_{z}v\|_{L^{\infty}}<\infty$ and $v(0,t)=0$ for all $t\in[0,T_{1}]$. Let $F_{i}:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\overline{\Omega}\times[0,T_{1}]\to\mathbb{R}$, $i=1,\dots,n$, be Lipschitz in the first two arguments uniformly on $\overline{\Omega}\times[0,T_{1}]$, i.e.
[…]
and bounded:
$\|F_{i}\|_{\infty}:=\sup_{(\mathbf{Y},\mathbf{C},z,t)}|F_{i}(\mathbf{Y},\mathbf{C})|<\infty.$
Assume $\mathbf{C}\in C(\overline{\Omega}\times[0,T_{1}];\mathbb{R}^{n})$ is given, and $\varphi_{i}\in C(\overline{\Omega})$ for $i=1,\dots,n$.
Then the system
$\left\{\begin{array}[]{ll}\partial_{t}Y_{i}-v(1,t)\,\partial_{z}Y_{i}=F_{i}(\mathbf{Y}(z,t),\mathbf{C}(z,t),z,t),&0<z<1,\;t>0,\\[3.00003pt] Y_{i}(z,0)=\varphi_{i}(z),&0\leq z\leq 1,\end{array}\right$. (13)
admits a unique mild (weak) solution $\mathbf{Y}\in C(\overline{\Omega}\times[0,T_{1}];\mathbb{R}^{n})$, and
$\|\mathbf{Y}\|_{\infty}\leq e^{LT_{1}}\big(\|\boldsymbol{\varphi}\|_{\infty}+T_{1}\|\mathbf{F}\|_{\infty}\big).$ (14)
If in addition each $\varphi_{i}\in C^{1}(\overline{\Omega})$ and $F_{i}$ are $C^{1}$ in $z$ with bounded derivatives, then the weak solution is classical, i.e
$\mathbf{Y}\in C^{1}(\overline{\Omega}\times[0,T_{1}];\mathbb{R}^{n}),$
and the the following estimate holds:
$\|\mathbf{Y}\|_{\infty}+\|\partial_{z}\mathbf{Y}\|_{\infty}\leq e^{LT_{1}}\big(\|\boldsymbol{\varphi}\|_{C^{1}}+T_{1}\|\mathbf{F}\|_{\infty}\big)+T_{1}e^{LT_{1}}\|\partial_{z}\mathbf{F}\|_{\infty}.$ (15)
Moreover, if $\varphi_{i}(z)\geq 0$ for all $z$ and $F_{i}(\mathbf{Y},\mathbf{C})\geq 0$ whenever $\mathbf{Y}\geq 0$, then
$Y_{i}(z,t)\geq 0,\qquad 0\leq z\leq 1,\;t\in[0,T_{1}],\;i=1,\dots,n.$
Proof. Since $v(1,t)$ depends only on $t$, we define the characteristic curve
$\frac{dZ}{ds}=-v(1,s),\qquad Z(0;z,t)=z.$
Because $v(1,\cdot)$ is continuous and bounded, there exists a unique absolutely continuous solution:
$Z(s;z,t)=z-\int_{s}^{t}v(1,\tau)\,d\tau.$
For fixed $(z,t)$ and each $i$, let $\zeta_{i}(s)=Y_{i}(Z(s;z,t),s)$. Then
$\frac{d\zeta_{i}}{ds}=\partial_{t}Y_{i}(Z(s),s)+\partial_{z}Y_{i}(Z(s),s)\frac{dZ}{ds}=\partial_{t}Y_{i}-v(1,s)\partial_{z}Y_{i}=F_{i}(\mathbf{Y}(Z(s),s),\mathbf{C}(Z(s),s)).$
Integrating from $s=0$ to $s=t$ gives the integral formulation
$Y_{i}(z,t)=\varphi_{i}(Z(0;z,t))+\int_{0}^{t}F_{i}(\mathbf{Y}(Z(s;z,t),s),\mathbf{C}(Z(s;z,t),s))\,ds.$ (16)
Define $\mathcal{T}:C(\overline{\Omega}\times[0,T_{1}];\mathbb{R}^{n})\to C(\overline{\Omega}\times[0,T_{1}];\mathbb{R}^{n})$ by
$(\mathcal{T}\mathbf{Y})_{i}(z,t):=\varphi_{i}(Z(0;z,t))+\int_{0}^{t}F_{i}(\mathbf{Y}(Z(s;z,t),s),\mathbf{C}(Z(s;z,t),s))\,ds.$
For $\mathbf{Y},\mathbf{\tilde{Y}}\in C$,
$|(\mathcal{T}\mathbf{Y})_{i}-(\mathcal{T}\mathbf{\tilde{Y}})_{i}|\leq\int_{0}^{t}L|\mathbf{Y}(Z(s),s)-\mathbf{\tilde{Y}}(Z(s),s)|\,ds\leq Lt\|\mathbf{Y}-\mathbf{\tilde{Y}}\|_{\infty}$.
Hence $\|\mathcal{T}\mathbf{Y}-\mathcal{T}\mathbf{\tilde{Y}}\|_{\infty}\leq LT_{1}\|\mathbf{Y}-\mathbf{\tilde{Y}}\|_{\infty}$. For $LT_{1}<1$, $\mathcal{T}$ is a contraction, so by the Banach fixed-point theorem, there exists a unique solution $\mathbf{Y}$ to (16) on $0,T_{1}]$. This $\mathbf{Y}$ is continuous and is the unique mild (weak) solution of (13). The $L^{\infty}$ estimate are obtained in the following way. From (16), one easily get
$|Y_{i}(z,t)|\leq\|\varphi_{i}\|_{\infty}+\int_{0}^{t}|F_{i}(\mathbf{Y}(Z(s),s),\mathbf{C}(Z(s),s))|\,ds.$
Since $|F_{i}(\mathbf{Y},\mathbf{C})|\leq L|\mathbf{Y}|+\|F_{i}\|_{\infty}$, we have
$|Y_{i}(z,t)|\leq\|\varphi_{i}\|_{\infty}+L\int_{0}^{t}|\mathbf{Y}(Z(s),s)|\,ds+t\|F_{i}\|_{\infty}.$
Let $\psi(t)=\max_{i}\sup_{z\in[0,1]}|Y_{i}(z,t)|$. Then
$\psi(t)\leq\|\boldsymbol{\varphi}\|_{\infty}+t\|\mathbf{F}\|_{\infty}+L\int_{0}^{t}\psi(s)\,ds.$
By Gronwall’s inequality,
$\psi(t)\leq e^{Lt}\big(\|\boldsymbol{\varphi}\|_{\infty}+t\|\mathbf{F}\|_{\infty}\big),$
and setting $t=T_{1}$ gives (14). To ensure the existence of a classical solution we assume now that $\varphi_{i}\in C^{1}(\overline{\Omega})$ and $F_{i}$ are $C^{1}$ in $z$ with bounded derivatives. Differentiating (16) with respect to $z$ gives
$\partial_{z}Y_{i}(z,t)=\varphi_{i}^{\prime}\big(Z(0;z,t)\big)+\int_{0}^{t}\partial_{z}F_{i}(\mathbf{Y}(Z(s),s),\mathbf{C}(Z(s),s))\,ds.$
Hence $\partial_{z}Y_{i}$ exists and is continuous, so $\mathbf{Y}\in C^{1}$. Moreover,
$|\partial_{z}Y_{i}(z,t)|\leq\|\varphi_{i}^{\prime}\|_{\infty}+t\|\partial_{z}F_{i}\|_{\infty}.$
Combining with the previous $L^{\infty}$ estimate and using again Gronwall’s lemma yields
$\|\mathbf{Y}\|_{\infty}+\|\partial_{z}\mathbf{Y}\|_{\infty}\leq e^{LT_{1}}\big(\|\boldsymbol{\varphi}\|_{C^{1}}+T_{1}\|\mathbf{F}\|_{\infty}\big)+T_{1}e^{LT_{1}}\|\partial_{z}\mathbf{F}\|_{\infty},$
which is (15). If $\varphi_{i}(z)\geq 0$ and $F_{i}(\mathbf{Y},\mathbf{C})\geq 0$ for $\mathbf{Y}\geq 0$, then from (16)
$Y_{i}(z,t)=\varphi_{i}(Z(0;z,t))+\int_{0}^{t}F_{i}(\mathbf{Y}(Z(s;z,t),s),\mathbf{C}(Z(s;z,t),s))\,ds\geq 0,$
so the solution remains nonnegative. This completes the proof. $\square$
For the local existence we will need the following assumptions in addition to those admititted in Lemmas 3.1 and 3.2.
(H1) For the initial and boundary and boundary data we assume the follwing
  • 1.
    $\boldsymbol{\varphi}=(\varphi_{1},\dots,\varphi_{n})\in C([0,1];\mathbb{R}^{n})$;
  • 2.
    $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{m})\in W^{2-2/p,p}(\Omega;\mathbb{R}^{m})$; $\boldsymbol{\Psi}=(\psi_{1},\dots,\psi_{m})\in W^{1,p}(0,T_{1};\mathbb{R}^{m})$ with the compatibility $\theta(1)=\Psi(0)$.
  • 3.
    $R_{0}>0$ is the initial domain size.
(H2) For coefficients and the reaction terms we consider the following regularity assumptions
  • 1.
    $v\in C^{1}(\overline{\Omega}\times[0,T_{1}])$ and $\|v\|_{\infty}+\|\partial_{z}v\|_{\infty}<\infty$.
  • 2.
    The reaction maps $\boldsymbol{F}=(F_{1},\dots,F_{n})$ and $\boldsymbol{H}=(H_{1},\dots,H_{m})$ are sufficiently regular: $\boldsymbol{F}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ and $\boldsymbol{H}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{m}$ are globally Lipschitz (in their vector arguments) and bounded on bounded sets. Moreover for the classical-regularity part assume $\boldsymbol{F}$ and $\boldsymbol{H}$ are $C^{1}$ in their arguments with bounded derivatives on the relevant range.
Next we give the existence and uniqueness theorem for the full coupled system studied in the project: The proof uses the two primary lemmas above and a contraction argument on a suitable product space.
Theorem 3.3
For $T_{1}>0$ and under the assumptions made above there exists $T^{\ast}\in(0,T_{1}]$ and a unique solution triple
$(\boldsymbol{Y},\boldsymbol{C},v,R)\quad\text{on }\overline{\Omega}\times[0,T^{\ast}]$
such that
  • 1.
    $\boldsymbol{C}\in W^{2,1}_{p}\big(\Omega\times(0,T^{\ast})\big)^{m}$ and satisfies the parabolic subsystem in the classical sense given by Lemma 3.1;
  • 2.
    $\boldsymbol{Y}\in C^{1}(\overline{\Omega}\times[0,T^{\ast}])^{n}$ and satisfies equation (4);
  • 3.
    $v\in C^{1}(\Omega\times[0,T^{*}])$ and $R\in C^{1}([0,T^{\ast}])$ satsify respectively (6) and (7) where $v(1,t)$ is the trace of the velocity at $z=1$ computed from $\boldsymbol{Y}$.
Moreover there exist constants $K_{1},K_{2}>0$ depending only on the data norms $\|\varphi\|_{C^{1}},\|\theta\|_{W^{2-2/p,p}},\|\psi\|_{W^{1,p}},R_{0}$, the Lipschitz constants, and $p$, such that for $0\leq t\leq T^{\ast}$:
$\|\boldsymbol{C}\|_{W^{2,1}_{p}((\Omega\times(0,t))}+\|\boldsymbol{Y}\|_{C^{1}(\overline{\Omega}\times[0,T^{*}])}+\|R\|_{C^{1}([0,T^{*}])}$ (17)
$\qquad\leq K_{1}\exp\!\big(K_{2}T^{*}\big)\Big(1+\|\varphi\|_{C^{1}}+\|\theta\|_{W^{2-2/p,p}}+\|\psi\|_{W^{1,p}}+R_{0}\Big).$
Finally this unique solution depends continuously on the data, and can be extended as long as the norms controlling the estimates remain finite.
Proof. The proof relies again on a fixed point argument on a small time interval. We split the construction into maps solved by the two given lemmas, $v$ and for $R$. By fixing $T>0$ to be chosen and we define the Banach space
$\mathcal{X}_{T}:=X_{Y}\times X_{R},\qquad X_{Y}:=\{\boldsymbol{Y}\in C^{1}(\overline{\Omega}\times[0,T])^{n}:\|\boldsymbol{Y}\|_{C^{1}}\leq M^{*}\},$
$X_{R}:=\{R\in C^{1}([0,T]):\|R\|_{C^{1}}\leq M^{*},\;R(0)=R_{0}\},$
with norm $\|(\boldsymbol{Y},R)\|_{\mathcal{X}_{T}}:=\|Y\|_{C^{1}}+\|R\|_{C^{1}}$. The constant $M^{*}>0$ can be chosen large enough (depending on the data) so the map we construct maps the closed ball $\mathcal{B}_{M}\subset\mathcal{X}_{T}$ into itself. By choosing arbitrary $(\boldsymbol{Y},R)\in\mathcal{B}_{M}$ we recall the parabolic right-hand side and transport coefficient
$H_{j}(z,t):=R^{2}(t)h_{j}\big(\boldsymbol{Y}(z,t),\boldsymbol{C}(z,t)\big),$
$\tilde{a}(t):=v(1,t)=R^{2}(t)\int_{0}^{1}g(\boldsymbol{Y}(\xi,t),\boldsymbol{C}(\xi,t))\,d\xi.$
Note that at this stage $\boldsymbol{C}$ is unknown inside the definition of $H_{j}$; we apply the parabolic lemma 3.1 in its mild nonlinear form: the lemma assumes the source depends Lipschitz continuously on the unknown $\boldsymbol{C}$ and gives a unique solution $\boldsymbol{C}\in W^{2,1}_{p}$ together with the estimate (for some constant $C_{P}$ depending on $p$, data and the norms of $Y,R$)
$\|\boldsymbol{C}\|_{W^{2,1}_{p}(\Omega\times(0,T))}\leq C_{P}\Big(\|\boldsymbol{\theta}\|_{W^{2-2/p,p}}+\|\boldsymbol{H}(\cdot,\cdot,0)\|_{L^{p}}+\|\boldsymbol{\Psi}\|_{W^{1,p}}\Big),$ (18)
where the dependence on $\boldsymbol{Y},R$ is via $\|\boldsymbol{Y}\|_{C^{1}}$ and $\|R\|_{C^{1}}$. By applying again the parabolic lemma 3.1 in its mild nonlinear form and by Sobolev embedding (since $p>3/2$) we also obtain
$\|\boldsymbol{C}\|_{C(\overline{\Omega}\times[0,T])}\leq C_{emb}\,\|\boldsymbol{C}\|_{W^{2,1}_{p}((0,1)\times(0,T))}$ (19)
for some $C_{emb}>0.$ Thus the parabolic lemma 3.1 furnishes a mapping
$\mathcal{P}:\mathcal{B}_{M}\to W^{2,1}_{p}((0,1)\times(0,T))^{m},\quad(\boldsymbol{Y},R)\mapsto\boldsymbol{C}.$
Using the $\boldsymbol{C}$ just obtained and the given $\boldsymbol{Y}$, we define
$v(z,t):=R^{2}(t)\int_{0}^{z}g\big(\boldsymbol{Y}(\xi,t),\boldsymbol{C}(\xi,t)\big)\,d\xi,$ (20a)
$v(1,t)=R^{2}(t)\int_{0}^{1}g\big(\boldsymbol{Y}(\xi,t),\boldsymbol{C}(\xi,t)\big)\,d\xi.$ (20b)
Because $g$ is Lipschitz and bounded and $\boldsymbol{Y},\boldsymbol{C},R$ are continuous, the integral defines $v\in C^{1}(\overline{\Omega}\times[0,T])$. Moreover, by the uniform bounds on $\boldsymbol{Y},\boldsymbol{C},R$ (obtained from the choice of $\mathcal{B}_{M}$ and (19)), the map $(\boldsymbol{Y},R)\mapsto v$ is Lipschitz into $C^{1}$ with a Lipschitz constant that is proportional to $T$ (because of the time integral structure when comparing two such $v$’s).
With the computed $v(1,t)$ and the right-hand side
$F_{i}(z,t):=R^{2}(t)f_{i}\big(\boldsymbol{Y}(z,t),\boldsymbol{C}(z,t)\big),$
apply the hyperbolic Lemma 3.2 to get the unique classical solution $\widetilde{Y}\in C^{1}(\overline{\Omega}\times[0,T])^{n}$ of
$\partial_{t}\widetilde{Y}_{i}-zv(1,t)\partial_{z}\widetilde{Y}_{i}=F_{i}(\cdot),\qquad\widetilde{Y}_{i}(\cdot,0)=\varphi_{i}.$
From Lemma 3.2 one can prove that there exists some constant $C_{H}$ depending on the Lipschitz constant of $f$, the data and the bound $M$ such that:
$\|\widetilde{Y}\|_{C^{1}(\overline{\Omega}\times[0,T])}\leq C_{H}\exp(C_{H}T)\Big(\|\varphi\|_{C^{1}}+T\|F\|_{C(\overline{\Omega}\times[0,T])}\Big).$ (21)
With these estimates above at hand, we now update $R$ by the ordinary differential equation. Thus, we define $\widetilde{v}(1,t)$ from (20) and solve the ODE
$\dot{\widetilde{R}}(t)=\widetilde{R}^{2}(t)\,\widetilde{v}(1,t)-\lambda\widetilde{R}^{4}(t),\qquad\widetilde{R}(0)=R_{0}.$
The right-hand side is locally Lipschitz in $\widetilde{R}$ and continuous in $t$, so the ODE has a unique $C^{1}$-solution on $[0,T]$. Moreover, using boundedness of $\widetilde{v}(1,t)$ and standard ODE estimates we get
$\|\widetilde{R}\|_{C^{1}([0,T])}\leq C_{R}\big(1+\|v(1,\cdot)\|_{C([0,T])}\big),$ (22)
with $C_{R}$ depending on $R_{0},\lambda$ and $T$.
We are now in position to define a solution operator that takes into account all the estimates established so far to deduce the fixed point result.
Let us introduce the operator
[…]
where $\widetilde{\boldsymbol{Y}}$ and $\widetilde{R}$ are produced above. The previous estimates show that for suitable $M$ large enough (depending on the data) and for $T$ sufficiently small the image $\mathcal{G}(\mathcal{B}_{M})\subset\mathcal{B}_{M}$. Indeed the right-hand sides of (18), (21), and (22) are bounded by constants depending on $M$ and the data. If we choose $M$ larger than those constants and then pick $T$ small enough so exponentials and $T$ factors still do not break the bounds.
It remains to get the contraction property. Indeed, let $(\boldsymbol{Y}^{1},R^{1}),(\boldsymbol{Y}^{2},R^{2})\in\mathcal{B}_{M}$ and denote the corresponding objects by $\boldsymbol{C}^{1},\boldsymbol{C}^{2}$, $v^{1},v^{2}$, $\widetilde{\boldsymbol{Y}}^{1},\widetilde{\boldsymbol{Y}}^{2}$, $\widetilde{R}^{1},\widetilde{R}^{2}$. We estimate the difference $\delta:=\|(\boldsymbol{Y}^{1},R^{1})-(\boldsymbol{Y}^{2},R^{2})\|_{\mathcal{X}_{T}}$.
From the Lemma 3.1 and since $\boldsymbol{H}$ is Lipschitz
$\|\boldsymbol{C}^{1}-\boldsymbol{C}^{2}\|_{W^{2,1}_{p}}\leq L_{C}(M)\big(\|\boldsymbol{Y}^{1}-\boldsymbol{Y}^{2}\|_{C(\overline{\Omega}\times[0,T])}+\|R^{1}-R^{2}\|_{C([0,T])}\big),$ (23)
hence by embedding,
$\|\boldsymbol{C}^{1}-\boldsymbol{C}^{2}\|_{C(\overline{\Omega}\times[0,T])}\leq L_{C,emb}(M)\,\delta.$
Then using the definition of $v$ and since and $g$ is also Lipschitz we have
$\|v^{1}-v^{2}\|_{C(\overline{\Omega}\times[0,T])}\leq T\cdot L_{v}(M)\,\delta,$
where the factor $T$ appears after integrating the difference of $g(\boldsymbol{Y},\boldsymbol{C})$ in time. Thus for small $T$ the mapping to the velocity is contractive in the sense that its Lipschitz constant contains a factor $T$.
Next, compare the Duhamel representations of $\widetilde{\boldsymbol{Y}}^{1}$ and $\widetilde{\boldsymbol{Y}}^{2}$ and use Lipschitzness of $\boldsymbol{F}$ and the estimate for $\|v^{1}-v^{2}\|$. One finds
$\|\widetilde{\boldsymbol{Y}}^{1}-\widetilde{\boldsymbol{Y}}^{2}\|_{C^{1}}\leq T\cdot L_{Y}(M)\,\delta,$ (24)
again with a small factor $T$ (the transport Duhamel integral gives explicitly a factor $t$ which is controlled by $T$).
Using the Gronwall lemma and the standard ODE theory we get that
$\|\widetilde{R}^{1}-\widetilde{R}^{2}\|_{C^{1}}\leq T\cdot L_{R}(M)\,\delta.$ (25)
Combining (24)–(25) we obtain, for a constant $L_{\mathcal{G}}(M)$,
$\|\mathcal{G}(\boldsymbol{Y}^{1},R^{1})-\mathcal{G}(\boldsymbol{Y}^{2},R^{2})\|_{\mathcal{X}_{T}}\leq L_{\mathcal{G}}(M)\,T\,\delta.$
Hence choosing $T$ so small that $L_{\mathcal{G}}(M)T<1$ we see that $\mathcal{G}$ is a contraction on $\mathcal{B}_{M}$.
By Banach fixed point theorem there exists a unique fixed point $(\boldsymbol{Y},R)\in\mathcal{B}_{M}$ with $\mathcal{G}(\boldsymbol{Y},R)=(\boldsymbol{Y},R)$. The parabolic solve with this $(\boldsymbol{Y},R)$ gives the corresponding $\boldsymbol{C}$. Thus the triple $(\boldsymbol{C},\boldsymbol{Y},R)$ is a solution of the coupled system on $[0,T]$. Uniqueness in the class $\boldsymbol{C}\in W^{2,1}_{p},\ \boldsymbol{Y}\in C^{1},\ R\in C^{1}$ follows from the contraction argument.
Combining the estimates from the parabolic Lemma (18), the transport Lemma (21), and the ODE estimate (22)yield bounds of the form
$\|\boldsymbol{C}\|_{W^{2,1}_{p}}\leq A_{1}+A_{2}\|\boldsymbol{Y}\|_{C^{1}}+A_{3}\|R\|_{C^{1}},\qquad\|\boldsymbol{Y}\|_{C^{1}}\leq B_{1}+B_{2}t\sup_{[0,t]}\|\boldsymbol{F}\|_{C},$
$\|R\|_{C^{1}}\leq N_{1}+N_{2}\|v(1,\cdot)\|_{C([0,t])},$
for constants depending on the data and Lipschitz constants. Iterating these inequalities and using Gronwall we obtain the stated exponential-type bound (17) (with constants $K_{1},K_{2}$ depending only on the data and the local ball radius $M$). This is the required a priori estimate on the local interval.
Finally, standard continuation arguments apply: the solution can be extended past $T^{\ast}$ as long as the norms appearing in $\|\boldsymbol{C}\|_{W^{2,1}_{p}}+\|\boldsymbol{Y}\|_{C^{1}}+\|R\|_{C^{1}}$ remain finite. This finishes the proof. $\square$
4 Global existence via invariance regions and energy estimates
This section is dedicated to extend the results iobtained in the previous sections for any $t\geq 0.$ To achieve this we need the following assumptions which extend the previous ones. Let us define $a(z,t)=zv(z,t).$ We have the following theorem
Theorem 4.1
Let $T_{1}>0$ and consider the coupled system (4) –(7) on $(z,t)\in\overline{\Omega}\times[0,T_{1}].$
Assume the following structural hypotheses hold.
  1. 1.
    $a\in C(\overline{\Omega}\times[0,T_{1}]),$ $a(1,t)=a(0,t)=0$ and there exists $M_{a}>0$ such that $\|a\|_{L^{\infty}}\leq M_{a}$.
  2. 2.
    (Diffusion) $D_{j}>0$ for all $j$.
  3. 3.
    (Lipschitzness) $f=(f_{i})_{i=1}^{n}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ and $h=(h_{j})_{j=1}^{m}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{m}$ are locally Lipschitz and satisfy the quasi-positivity property:
    $f_{i}(\boldsymbol{Y},\boldsymbol{C})\geq 0\ \text{ whenever }\boldsymbol{Y}\geq\boldsymbol{0},\boldsymbol{C}\geq\boldsymbol{0},$
    $h_{j}(\boldsymbol{Y},\boldsymbol{C})\geq\boldsymbol{0}\ \text{ whenever }\boldsymbol{Y}\geq\boldsymbol{0},\boldsymbol{C}\geq\boldsymbol{0}.$
  4. 4.
    (Dissipative energy inequality) There exist constants $\alpha>0$, $\beta\geq 0$ and $M_{0}\geq 0$ such that for all $Y\in\mathbb{R}^{n}$, $C\in\mathbb{R}^{m}$ and all $t\geq 0$,
    $\sum_{i=1}^{n}\mu_{i}f_{i}(\boldsymbol{Y},\boldsymbol{C})Y_{i}+\sum_{j=1}^{m}\nu_{j}h_{j}(\boldsymbol{Y},\boldsymbol{C})C_{j}\leq-\alpha\big(\|\boldsymbol{Y}\|_{2}^{2}+\|\boldsymbol{C}\|_{2}^{2}\big)+\beta+M_{0},$ (26)
    for some fixed positive weights $\mu_{i},\nu_{j}>0$ (one may take $\mu_{i}=\nu_{j}=1$).
Then the unique local classical solution given by the local existence theorem 3.3 extends to a global classical solution on $\overline{\Omega}\times[0,\infty)$ with
$\boldsymbol{Y}\in C^{1}(\overline{\Omega}\times[0,\infty))^{n},~\boldsymbol{C}\in W^{2,1}_{p,\mathrm{loc}}(\overline{\Omega}\times[0,\infty))^{m},~v\in C^{1}(\overline{\Omega}\times[0,\infty))^{n},R\in C^{1}([0,\infty)).$
Moreover, the solution remains nonnegative and satisfies the uniform in time energy bound
$\|\boldsymbol{Y}(\cdot,t)\|_{L^{2}}^{2}+\|\boldsymbol{C}(\cdot,t)\|_{L^{2}}^{2}+\|v(\cdot,t)\|_{L^{2}}^{2}+R^{2}(t)\leq C_{\infty}\qquad\text{for all }t\geq 0,$ (27)
for some constant $C_{\infty}$ depending only on the initial data and model constants.
This inequality (27) is a dissipation condition which ensures that reaction terms do not allow arbitrary growth and provide a restoring effect for large norms. To prove this system we use the local existence result and the continuation criterion then derive a priori energy estimates which are uniform on finite time intervals and in fact globally bounded. We conclude the proof by using these estimates to prevent the norms that control continuation from blowing up; hence extend the solution globally.
Proof. By the local well-posedness result in Theorem 3.3) we only need to obtain a priori bounds preventing blow-up of the estimates obtained earlier. In addition, since each $f_{i}$ and $h_{j}$ are quasi-positive (hypothesis 3)) and the transport and diffusion operators preserve sign under the given boundary conditions, standard comparison-principle arguments yield
$Y_{i}(z,t)\geq 0,\qquad C_{j}(z,t)\geq 0\qquad\text{for all }(z,t)\in[0,1]\times[0,T_{\mathrm{loc}}].$
Indeed, if at some time a component would become negative, the quasi-positivity of the source forbids a decrease through the source term; transport and diffusion alone preserve nonnegativity when initial or boundary data are nonnegative. From the local solution this property holds on $0,T_{\mathrm{loc}}]$. Thus the region $\{\boldsymbol{Y}\geq 0,\boldsymbol{C}\geq 0\}$ is invariant.
To get the energy estimates we multiply each biomass equation (4) by $\mu_{i}Y_{i}$ and integrate over $z\in(0,1)$; multiply each substrate equation (5a) by $\nu_{j}C_{j}$ and integrate, then sum the results (weights $\mu_{i},\nu_{j}>0$ are those appearing in (26)). Using integration by parts for the diffusion terms we obtain
$\frac{d}{dt}\Big\{\tfrac{1}{2}\|\boldsymbol{Y}(\cdot,t)\|_{L^{2}_{\mu}}^{2}+\tfrac{1}{2}\|\boldsymbol{C}(\cdot,t)\|_{L^{2}_{\nu}}^{2}\Big\}+\;\sum_{j=1}^{m}\nu_{j}D_{j}\int_{0}^{1}|\partial_{z}C_{j}|^{2}\,dz$
$=\;R^{2}(t)\int_{0}^{1}\Big(\sum_{i=1}^{n}\mu_{i}f_{i}(\boldsymbol{Y},\boldsymbol{C})Y_{i}+\sum_{j=1}^{m}\nu_{j}h_{j}(\boldsymbol{Y},\boldsymbol{C})C_{j}\Big)\,dz$
$\qquad+\int_{0}^{1}\Big(zv(1,t)\partial_{z}C_{j}C_{j}+zv(1,t)\partial_{z}Y_{i}Y_{i}\Big)dz,$ (28)
where $\|Y\|_{L^{2}_{\mu}}^{2}=\sum_{i}\mu_{i}\int_{0}^{1}Y_{i}^{2}\,dz$ and similarly for $C$. Using the hhypothesis 1, the last two integral vanish. For generality we assume boundary contributions are controlled (they are either zero under common boundary conditions or can be estimated by the $L^{2}$-norms of $\boldsymbol{Y},\boldsymbol{C}$).
Next, we apply (26) pointwise in $(z,t)$ and integrate over $z$ to estimate the remainder of right-hand side of (4):
$R^{2}(t)\int_{0}^{1}\Big(\sum_{i}\mu_{i}f_{i}Y_{i}+\sum_{j}\nu_{j}h_{j}C_{j}\Big)\,dz\leq R^{2}(t)\int_{0}^{1}\big(-\alpha(\|Y\|_{2}^{2}+\|C\|_{2}^{2})+\beta+M_{0}\big)\,dz.$
Because $\int_{0}^{1}(\|\boldsymbol{Y}\|_{2}^{2}+\|\boldsymbol{C}\|_{2}^{2})\,dz=\|\boldsymbol{Y}\|_{L^{2}}^{2}+\|\boldsymbol{C}\|_{L^{2}}^{2}$, we get (omitting the boundary term for clarity)
$\frac{d}{dt}\mathcal{E}(t)+\kappa\|\partial_{z}\boldsymbol{C}(\cdot,t)\|_{L^{2}}^{2}\leq-\alpha R^{2}(t)\big(\|\boldsymbol{Y}(\cdot,t)\|_{L^{2}}^{2}+\|\boldsymbol{C}(\cdot,t)\|_{L^{2}}^{2}\big)+R^{2}(t)(\beta+M_{0}),$ (29)
where we have set the energy
$\mathcal{E}(t):=\tfrac{1}{2}\big(\|\boldsymbol{Y}(\cdot,t)\|_{L^{2}_{\mu}}^{2}+\|\boldsymbol{C}(\cdot,t)\|_{L^{2}_{\nu}}^{2}\big)$
and $\kappa=\min_{j}\nu_{j}D_{j}>0$. To bound R we recall that it satisfies the ODE (7). One notices that the ODE contains a strong damping term $\lambda R^{4}$ which prevents runaway of $R$ for large values. More precisely if we consider the function $Q(t):=R^{2}(t)$, then we have
$\dot{Q}(t)=2R(t)\dot{R}(t)=2R^{3}(t)v(t)-2\lambda R^{5}(t)=2Q^{3/2}v(t)-2\lambda Q^{5/2}.$
From this we get a differential inequality that shows $R(t)$ cannot blow up in finite time because for large $Q$ the $2\lambda Q^{5/2}$ term dominates and forces a decrease. More directly, using standard comparison with the autonomous ODE $\dot{R}=-\tfrac{1}{2}\lambda R^{4}$ for sufficiently large $R$ and since $v$ is bounded, we further infer that there exists $R_{\max}$ depending on $R_{0},\lambda$ and $\|v\|_{\infty}$ such that
$R(t)\leq R_{\max}\qquad\text{for all }t\geq 0.$
In particular $R^{2}(t)$ is globally bounded. Notably there exists $M_{R}>0$ such that $R^{2}(t)\leq M_{R}$ for all $t\geq 0$ makes sense.
Using the bound $R^{2}(t)\leq M_{R}$ in (29) yields
$\frac{d}{dt}\mathcal{E}(t)\leq-\alpha M_{R}\big(\|\boldsymbol{Y}\|_{L^{2}}^{2}+\|\boldsymbol{C}\|_{L^{2}}^{2}\big)+M_{R}(\beta+M_{0}).$
Dropping the negative first term on the right gives the crude linear bound
$\frac{d}{dt}\mathcal{E}(t)\leq M_{R}(\beta+M_{0}).$
Integrating from 0 to $t$,
$\mathcal{E}(t)\leq\mathcal{E}(0)+M_{R}(\beta+M_{0})\,t.$
This alone gives at most linear growth. To obtain a uniform-in-time bound we use the full dissipative structure and rewrite (29) as
$\frac{d}{dt}\mathcal{E}(t)\leq-\alpha M_{R}\big(2\mathcal{E}(t)/C^{*}\big)+M_{R}(\beta+M_{0}),$
where $C^{*}>0$ is a constant comparing the weighted and usual $L^{2}$-norms: $\|\boldsymbol{Y}\|_{L^{2}_{\mu}}^{2}+\|\boldsymbol{C}\|_{L^{2}_{\nu}}^{2}\geq C^{*}(\|\boldsymbol{Y}\|_{L^{2}}^{2}+\|\boldsymbol{C}\|_{L^{2}}^{2})$. Thus
$\frac{d}{dt}\mathcal{E}(t)\leq-\gamma\mathcal{E}(t)+M_{R}(\beta+M_{0}),\qquad\text{with }\gamma=\frac{2\alpha M_{R}}{C^{*}}>0.$
By the standard linear Gronwall inequality for this linear dissipative ODE we obtain
$\mathcal{E}(t)\leq e^{-\gamma t}\mathcal{E}(0)+\frac{M_{R}(\beta+M_{0})}{\gamma}\leq\mathcal{E}(0)+\frac{M_{R}(\beta+M_{0})}{\gamma}=:E_{\infty}.$
Hence $\mathcal{E}(t)$ is uniformly bounded for all $t\geq 0$.
Consequently there exists $C_{\infty}>0$ so that
$\|\boldsymbol{Y}(\cdot,t)\|_{L^{2}}^{2}+\|\boldsymbol{C}(\cdot,t)\|_{L^{2}}^{2}\leq C_{\infty},\qquad\forall t\geq 0,$
which, together with the bound for $R(t)$, yields (27). Notice the velocity bound can also be deduced by using the fact that $g$ is sum of terms from $f_{i}$ (which are bounded) and boundeness of $R.$
We now upgrade the $L^{2}$-bounds to the norms used in the continuation criterion. For the parabolic variables $\boldsymbol{C}$: since the right-hand side $R^{2}(t)h_{j}(\boldsymbol{Y},\boldsymbol{C})$ is uniformly bounded in $L^{2}(\Omega)$ by the $L^{2}$-bounds on $\boldsymbol{Y},\boldsymbol{C}$ and the uniform bound on $R$, classical $L^{2}$-parabolic regularity estimates yield that for any finite time $T>0$,
$\|\boldsymbol{C}\|_{W^{2,1}_{p}(\Omega\times(0,T))}\leq C_{G}(T),$
for some finite constant $C_{G}(T)$ depending only on the initial data and the bounds obtained above (and on $T$). In particular for each finite $T$ the $\boldsymbol{C}$-norms used in the continuation criterion remain finite.
For the transport variables $\boldsymbol{Y}$: the transport equation admits uniform $L^{\infty}$–bounds because the right-hand side $R^{2}f_{i}(\boldsymbol{Y},\boldsymbol{C})$ is controlled by the $L^{2}$-bounds on $Y,C$ and the Lipschitz structure of $f_{i}$ together with invariant region arguments. More precisely, using the Duhamel integral formula along characteristics and Gronwall-type arguments, one deduces for each finite $T$ a bound
$\|\boldsymbol{Y}\|_{C^{1}(\overline{\Omega}\times[0,T])}\leq C_{Y}(T),$
again depending only on the a priori quantities already bounded. Finally, $R(t)$ is bounded and continuous for all time as shown earlier. Therefore none of the norms that control the local continuation criterion blow up in finite time.
Since all continuation norms remain bounded for all finite times, the local solution can be extended step by step to a global solution on $0,\infty)$. The uniform-in-time a priori bounds (27) prevent finite-time blow-up and guarantee classical regularity persists for all times. $\square$